OpenStudy (anonymous):

The body mass index, I, can be used to determine an individuals risk for heart disease. An index less than 25 indicates low risk. I=700w/h^2 where w=weight in pounds, h=height in inches. Jerome is 68 inches tall. What weights will keep his body mass index between 18 and 31

OpenStudy (tad1):

18 = (700)(w/68^2 31 = (700)(w)/68^2 work out both equations, your answer w wil be an inequality between the two results

OpenStudy (anonymous):

The answer should look like this ____<w<_______ I think that it should be 18<W<31

OpenStudy (tad1):

no, w is the weight. You have to solve each of the equaitons I gave you. the form is right, but 18 and 31 are indices not weights.

OpenStudy (anonymous):

OOOOOOOOOHHHH ok :( I will try...

OpenStudy (saifoo.khan):

hey, i solved that.. msgd on fb!

OpenStudy (anonymous):

hey I seent you a message back.